OpenStudy (anonymous):

A rectangular prism is truncated removing 20% of its height. The original height of the prism was 100 inches and it has a length of 5 inches and a width of 10 inches. What is the volume of the truncated prism? Answer 2,000 in3 3,000 in3 4,000 in3 5,000 in3 for this one i got 3,000?

OpenStudy (anonymous):

the height was 100inches now to take off 20% of 100 we use: \[\LARGE 100(1-20\%)=\] \[\LARGE 100-100\cdot {20\over 100 }=\] \[\LARGE 100-20=80\] so we have its new cordinates: height: H=80 inches Length 5 inches and width 10 inches now we have:  Although the picture doesn't fit with its dimensions, let's assume they're ok. :) and we have the formula for Volume: \[\LARGE V=l\cdot h\cdot w\] \[\LARGE V=5 \;\; inch \cdot 80 \;\; innch \cdot 10\;\;inch\] can you do it now ? :)